# A vectorial Darboux transformation for integrable matrix versions of the Fokas-Lenells equation

### v1

@@ 6.1.1 The case $n=1$
Then the general solution of the linear system is
$\chi_{1}=\Big{(}c_{1}+c_{2}\,\big{(}x-\mathrm{i}\,\frac{\gamma}{\alpha}\,t\big{)}\Big{)}\,e^{\frac{\mathrm{i}\,}{2}\,(\alpha\,x+\beta\,t)}\,,\qquad\chi_{2}=c_{0}\,e^{\frac{1}{2\gamma}\,x+\frac{\gamma}{2}\,t}\,e^{\mathrm{i}\,\,(\alpha\,x+\beta\,t)}\,$,
with complex constants $c_{0},c_{1},c_{2}$ and $\beta=2\|A\|^{2}-\alpha^{-1}$. If the real discriminant in the expression for $\gamma$ is non-positive, then $\gamma$ is imaginary and the constraint resulting from the Lyapunov equation enforces $c_{2}=0$, which takes us back to the case covered by Example 6.4(ii). In the remaining case, where $\alpha^{-1}>\|A\|^{2}$, we have $\mathrm{Re}(\gamma)\neq 0$. The Lyapunov equation then has the unique solution
$\Omega=\frac{1}{2\,\mathrm{Re}(\gamma)}\big{(}\frac{\mathrm{i}\,\gamma^{\ast}}{\alpha^{2}\|A\|^{4}}|\Xi|^{2}+|\chi_{1}|^{2}+|\chi_{2}|^{2}\big{)}\,$,
where
$\Xi=\|A\|\,\Big{(}c_{2}\,e^{\frac{\mathrm{i}\,}{2}\,(\alpha\,x+\beta\,t)}-\frac{1}{2}(\gamma^{-1}+\mathrm{i}\,\alpha)\,\chi_{1}\Big{)}\,.$
The generated solution of the two-component FL equation is now obtained by inserting the expressions for $\chi_{1},\chi_{2},\Xi,\Omega$ in (6.21). The solution is quasi-rational if $c_{0}=0$. Then it models a rogue wave, rationally localized in space and time. Fig. 6 displays an example.
@@
If $A_{1}=A_{2}$, we have $u_{1}=u_{2}$. This is not true if $c_{0}\neq 0$, in which case we have a superposition with a breather, see Fig. 7.
@@
The solution set of a polynomial matrix equation, like (6.13), is in general more intricate than for polynomial equations over a field. Here we take a look at an example in our context. Let
@@
$\Gamma=\left(\begin{array}[]{cc}\gamma_{1}&0\\ 0&\gamma_{2}\end{array}\right)\,$,
where $\gamma_{1},\gamma_{2}$ have the form (6.34). In this case,
@@
$\chi_{1}=e^{\frac{\mathrm{i}\,}{2}\,(\alpha\,x+\beta\,t)}\left(\begin{array}[]{c}c_{11}+c_{12}\,(x-\frac{\mathrm{i}\,}{\alpha}\gamma_{1}t)\\ c_{21}+c_{22}\,(x-\frac{\mathrm{i}\,}{\alpha}\gamma_{2}t)\end{array}\right)\,,\qquad\chi_{2}=e^{\mathrm{i}\,(\alpha\,x+\beta\,t)}\left(\begin{array}[]{c}c_{01}\,e^{\frac{1}{2}(\gamma_{1}^{-1}x+\gamma_{1}t)}\\ c_{02}\,e^{\frac{1}{2}(\gamma_{2}^{-1}x+\gamma_{2}t)}\end{array}\right)\,.$
@@
Appendix C presents some examples of superpositions of two solutions from the class in Example 6.2.
6.2 The generic case $\alpha_{1}\neq\alpha_{2}$

### v2

@@ 6.1.1 $n=1$
Then the general solution of (6.12) is
$\chi_{1}=(c_{1}+c_{2}\,\rho)\,e^{\frac{\mathrm{i}}{2}\,\varphi}\,,\qquad\rho=x-\mathrm{i}\,\frac{\gamma}{\alpha}\,t\,$,
with complex constants $c_{1},c_{2}$. If the real discriminant in the expression for $\gamma$ is non-positive, then $\gamma$ is imaginary and the constraint resulting from the Lyapunov equation enforces $c_{2}=0$, which takes us back to the case covered under (ii) in Example 6.7. In the remaining case, where $\alpha^{-1}>\|A\|^{2}$, we have $\mathrm{Re}(\gamma)\neq 0$. We obtain
$\Xi=-\frac{\|A\|}{2\gamma}\,\Big((1+\mathrm{i}\,\alpha\gamma)(c_{1}+c_{2}\,\rho)-2c_{2}\gamma\Big)\,e^{\frac{\mathrm{i}}{2}\,\varphi}\,$,
and the Lyapunov equation has the unique solution
$\Omega=\frac{1}{2\,\mathrm{Re}(\gamma)}\big(\frac{\mathrm{i}\,\gamma^{\ast}}{\alpha^{2}\|A\|^{4}}|\Xi|^{2}+|\chi_{1}|^{2}+|\chi_{2}|^{2}\big)\,.$
According to (6.21), the generated solution of the two-component FL equation is now given by
$u^{\prime}=e^{\mathrm{i}\,\varphi}\,\boldsymbol{U}_{0}\,\left(\begin{array}[]{c}\|A\|+(\alpha\,\|A\|^{2}\,\Omega)^{-1}\chi_{1}^{\ast}\,\Xi\\ (\alpha\,\|A\|^{2}\,\Omega)^{-1}\,\chi_{2}^{\ast}\,\Xi\end{array}\right)\,.$
This solution is quasi-rational if $c_{0}=0$ and then models a rogue wave, rationally localized in space and time and equivalent to a corresponding solution of the scalar FL equation. Fig. 7 displays an example.
@@
If $c_{0}\neq 0$, the solution is not equivalent to one of the scalar equation. In this case, we have a superposition with a breather, see Fig. 8 for an example. The double root case treated in this example has already been considered in [12].
@@
Appendix Appendix D: Superposition of two solutions from the class in Example 6.5 presents some examples of superpositions of two solutions from the class in Example 6.5.
6.1.2 $n=2$
If $\Gamma$ is diagonal and either satisfies the spectrum condition or has imaginary eigenvalues (which is one of the cases where the spectrum condition does not hold), then the corresponding $n=2$ solution is easily obtained by use of results in Section 6.1.1 and an application of the superposition rule, formulated in Remark 5.6. Next we address another case, where $\Gamma$ is diagonal and the spectrum condition is violated, and then the case, where $\Gamma$ is a Jordan block.
@@ 6.1.2 $n=2$
Let $\Gamma=\mathrm{diag}(\gamma,-\gamma^{\ast})$ with $\mathrm{Re}(\gamma)\neq 0$. Then $R=\mathrm{diag}(\pm r,\pm r^{\ast})$ with
$r=\frac{1}{2}\sqrt{\gamma^{-2}-2\mathrm{i}\,\alpha^{2}\beta\,\gamma^{-1}-\alpha^{2}}\,.$
Although the matrix equation for $R$ thus has four different matrix roots, it is plausible that already two of them are sufficient to obtain the complete solution of the linear system. Assuming $r\neq 0$, this is
@@
$\chi_{1}=e^{\frac{\mathrm{i}}{2}\varphi}\left(\begin{array}[]{c}c_{11}\,e^{\psi}+c_{12}\,e^{-\psi}\\ c_{21}\,e^{\psi^{\ast}}+c_{22}\,e^{-\psi^{\ast}}\end{array}\right)\,,\quad\chi_{2}=e^{\mathrm{i}\,\varphi}\left(\begin{array}[]{c}c_{01}\,e^{\frac{1}{2}(\gamma^{-1}x+\gamma\,t)}\\ c_{02}\,e^{-\frac{1}{2}(\gamma^{\ast-1}x+\gamma^{\ast}\,t)}\end{array}\right)\,$,
where
$\psi=r\,(x-\mathrm{i}\,\alpha^{-1}\gamma\,t)\,.$
We further obtain
$\Xi=\frac{\|A\|}{2}\,e^{\frac{\mathrm{i}}{2}\varphi}\left(\begin{array}[]{c}c_{11}\,(2r-\gamma^{-1}-\mathrm{i}\,\alpha)\,e^{\psi}-c_{12}\,(2r+\gamma^{-1}+\mathrm{i}\,\alpha)\,e^{-\psi}\\ c_{21}\,(2r^{\ast}+\gamma^{\ast-1}-\mathrm{i}\,\alpha)\,e^{\psi^{\ast}}-c_{22}\,(2r^{\ast}-\gamma^{\ast-1}+\mathrm{i}\,\alpha)\,e^{-\psi^{\ast}}\end{array}\right)\,.$
The Lyapunov equation determines the diagonal entries of $\Omega$,
$\Omega_{11}$ $=$ $\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\Xi\Xi^{\dagger}\,\Gamma^{\dagger}+\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger}\Big)_{11}$
$=$ $\frac{1}{2\mathrm{Re}(\gamma)}\Big(\big(\frac{\mathrm{i}\,\gamma^{\ast}}{4\alpha^{2}\|A\|^{2}}|2r-\gamma^{-1}-\mathrm{i}\,\alpha|^{2}+1\big)\,|c_{11}|^{2}e^{2\mathrm{Re}(\psi)}$
$-\big(\frac{\mathrm{i}\,\gamma^{\ast}}{4\alpha^{2}\|A\|^{2}}(2r-\gamma^{-1}-\mathrm{i}\,\alpha)\,(2r^{\ast}+\gamma^{\ast-1}-\mathrm{i}\,\alpha)-1\big)\,c_{11}c_{12}^{\ast}e^{2\mathrm{i}\,\mathrm{Im}(\psi)}$
$-\big(\frac{\mathrm{i}\,\gamma^{\ast}}{4\alpha^{2}\|A\|^{2}}(2r+\gamma^{-1}+\mathrm{i}\,\alpha)\,(2r^{\ast}-\gamma^{\ast-1}+\mathrm{i}\,\alpha)-1\big)\,c_{11}^{\ast}c_{12}e^{-2\mathrm{i}\,\mathrm{Im}(\psi)}$
$+\big(\frac{\mathrm{i}\,\gamma^{\ast}}{4\alpha^{2}\|A\|^{2}}|2r+\gamma^{-1}+\mathrm{i}\,\alpha|^{2}+1\big)\,|c_{12}|^{2}e^{-2\mathrm{Re}(\psi)}+|c_{01}|^{2}e^{\mathrm{Re}(\gamma^{-1}x+\gamma t)}\Big)\,$,
$\Omega_{22}$ $=$ $-\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\Xi\Xi^{\dagger}\,\Gamma^{\dagger}+\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger}\Big)_{22}$
$=$ $\frac{1}{2\mathrm{Re}(\gamma)}\Big(\frac{\mathrm{i}\,\gamma}{4\alpha^{2}\|A\|^{2}}\big(|2r+\gamma^{-1}+\mathrm{i}\,\alpha|^{2}-1\big)\,|c_{21}|^{2}e^{2\mathrm{Re}(\psi)}$
$-\big(\frac{\mathrm{i}\,\gamma}{4\alpha^{2}\|A\|^{2}}(2r+\gamma^{-1}+\mathrm{i}\,\alpha)\,(2r^{\ast}-\gamma^{\ast-1}+\mathrm{i}\,\alpha)+1\big)\,c_{21}^{\ast}c_{22}e^{2\mathrm{i}\,\mathrm{Im}(\psi)}$
$-\big(\frac{\mathrm{i}\,\gamma}{4\alpha^{2}\|A\|^{2}}(2r-\gamma^{-1}-\mathrm{i}\,\alpha)\,(2r^{\ast}+\gamma^{\ast-1}-\mathrm{i}\,\alpha)+1\big)\,c_{21}c_{22}^{\ast}e^{-2\mathrm{i}\,\mathrm{Im}(\psi)}$
$+\frac{\mathrm{i}\,\gamma}{4\alpha^{2}\|A\|^{2}}\big(|2r-\gamma^{-1}-\mathrm{i}\,\alpha|^{2}-1\big)\,|c_{22}|^{2}e^{-2\mathrm{Re}(\psi)}-|c_{02}|^{2}e^{-\mathrm{Re}(\gamma^{-1}x+\gamma t)}\Big)\,$,
and imposes the constraints
$\Big(\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\Xi\Xi^{\dagger}\,\Gamma^{\dagger}+\chi_{1}\chi_{1}^{\dagger}+\chi_{2}\chi_{2}^{\dagger}\Big)_{ij}=0\qquad\mbox{for}\quad i\neq j\,.$
The latter amount to
$(c_{11}c_{22}^{\ast}+c_{12}c_{21}^{\ast})\,\big(4\alpha^{2}\|A\|^{2}\gamma+\mathrm{i}\,(1+\mathrm{i}\,\alpha\gamma)^{2}\big)-2\mathrm{i}\,\gamma\,r\,(c_{11}c_{22}^{\ast}-c_{12}c_{21}^{\ast})\,(1+\mathrm{i}\,\alpha\gamma)+2c_{01}c_{02}^{\ast}\,\alpha^{2}\|A\|^{2}\gamma=0\,.$
The differential equations for $\Omega$ lead to
$\Omega_{12x}$ $=$ $\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\big(\Xi_{1x}-\frac{1}{2}\gamma^{-1}\Xi_{1}\big)\,\Xi_{2}^{\ast}\,,\qquad\Omega_{12t}=-\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\big(\alpha\|A\|^{3}\chi_{12}^{\ast}-\gamma\,\Xi_{2}^{\ast}\big)\,\Xi_{1}\,$,
$\Omega_{21x}$ $=$ $\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\big(\Xi_{2x}+\frac{1}{2}\gamma^{\ast-1}\Xi_{2}\big)\,\Xi_{1}^{\ast}\,,\qquad\Omega_{21t}=-\frac{\mathrm{i}}{\alpha^{2}\|A\|^{4}}\big(\alpha\|A\|^{3}\chi_{11}^{\ast}+\gamma^{\ast}\Xi_{1}^{\ast}\big)\,\Xi_{2}\,$,
which leads to
$\Omega_{12}$ $=$ $\frac{\mathrm{i}}{16\alpha^{2}\|A\|^{2}\gamma^{3}\,r}\Big(c_{11}c_{21}^{\ast}\big(1+\gamma\,(\mathrm{i}\,\alpha+2r)\big)\,\big(1+\gamma^{2}\,(\alpha^{2}+4r^{2})-4\gamma\,r\big)\,e^{2\psi}$
$-c_{12}c_{22}^{\ast}\big(1+\gamma\,(\mathrm{i}\,\alpha-2r)\big)\,\big(1+\gamma^{2}(\alpha^{2}+4r^{2})+4\gamma\,r\big)\,e^{-2\psi}\Big)$
$+\frac{\mathrm{i}\,x}{8\alpha^{2}\|A\|^{2}\gamma^{3}}\Big(c_{12}c_{21}^{\ast}\big(1+\gamma\,(\mathrm{i}\,\alpha+2r)\big)\,\big(1+\gamma^{2}\,(\alpha^{2}+4r^{2})+4\gamma\,r\big)$
$+c_{11}c_{22}^{\ast}\,\big(1+\gamma\,(\mathrm{i}\,\alpha-2r)\big)\,\big(1+\gamma^{2}(\alpha^{2}+4r^{2})-4\gamma\,r\big)\Big)$
$+\frac{\mathrm{i}\,t}{4\alpha^{2}\|A\|^{2}\gamma}\Big(c_{12}c_{21}^{\ast}\,\big(2\alpha\|A\|^{2}\,(1+\gamma\,(\mathrm{i}\,\alpha+2r)\big)+\gamma^{2}\,\big(\alpha^{2}-4r\,(\mathrm{i}\,\alpha+r)\big)-2\gamma\,(\mathrm{i}\,\alpha+2r)-1\big)$
$+c_{11}c_{22}^{\ast}\,\big(2\alpha\|A\|^{2}\,\big(1+\gamma\,(\mathrm{i}\,\alpha-2r)\big)+\gamma^{2}\,\big(\alpha^{2}+4r\,(\mathrm{i}\,\alpha-r)\big)-2\gamma\,(\mathrm{i}\,\alpha-2r)-1\big)\Big)+\tilde{c}_{12}\,$,
$\Omega_{21}$ $=$ $-\frac{\mathrm{i}}{16\alpha^{2}\|A\|^{2}\gamma^{\ast 3}\,r^{\ast}}\Big(c_{11}^{\ast}c_{21}\big(1-\gamma^{\ast}(\mathrm{i}\,\alpha+2r^{\ast})\big)\,\big(1+\gamma^{\ast 2}\,(\alpha^{2}+4r^{\ast 2})+4\gamma^{\ast}r^{\ast}\big)\,e^{2\psi^{\ast}}$
$-c_{12}^{\ast}c_{22}\big(1-\gamma^{\ast}(\mathrm{i}\,\alpha-2r^{\ast})\big)\,\big(1+\gamma^{\ast 2}(\alpha^{2}+4r^{\ast 2})-4\gamma^{\ast}r^{\ast}\big)\,e^{-2\psi^{\ast}}\Big)$
$+\frac{\mathrm{i}\,x}{8\alpha^{2}\|A\|^{2}\gamma^{\ast 3}}\Big(2\gamma^{\ast}r^{\ast}(c_{11}^{\ast}c_{22}-c_{12}^{\ast}c_{21})\,\big(\gamma^{\ast 2}(\alpha^{2}+4r^{\ast 2})-2\mathrm{i}\,\alpha\gamma^{\ast}+3\big)$
$+(c_{11}^{\ast}c_{22}+c_{12}^{\ast}c_{21})\,\big(4r^{\ast 2}(\mathrm{i}\,\alpha\gamma^{\ast 3}-3\gamma^{\ast 2})+(1+\alpha^{2}\gamma^{\ast 2})\,(\mathrm{i}\,\alpha\gamma^{\ast}-1)\big)\Big)$
$+\frac{\mathrm{i}\,t}{4\alpha^{2}\|A\|^{2}\gamma^{\ast}}\Big(4\gamma^{\ast}r^{\ast}(c_{11}^{\ast}c_{22}-c_{12}^{\ast}c_{21})\,\big(\alpha\,(\mathrm{i}\,\gamma^{\ast}+\|A\|^{2})-1\big)$
$+(c_{11}^{\ast}c_{22}+c_{12}^{\ast}c_{21})\,\big(1-\gamma^{\ast 2}(\alpha^{2}-4r^{\ast 2})+2\alpha\|A\|^{2}(\mathrm{i}\,\alpha\gamma^{\ast}-1)-2\mathrm{i}\,\alpha\gamma^{\ast}\big)\Big)+\tilde{c}_{21}\,$,
with complex constants $\tilde{c}_{12}$ and $\tilde{c}_{21}$. The remaining condition (4.5) becomes
$\gamma\,(\tilde{c}_{12}-\tilde{c}_{21}^{\ast})=c_{01}c_{02}^{\ast}+c_{11}c_{22}^{\ast}+c_{12}c_{21}^{\ast}\,.$
The resulting solution of the two-component FL equation is obtained by inserting the expressions for $\chi_{1}$, $\chi_{2}$, $\Xi$ and $\Omega$ in (6.21). In addition to the exponential dependence of the solution via $e^{\psi}$ and its complex conjugate, there is also a rational dependence on the variables $x$ and $t$. Fig. 9 displays an example from this class.
Figure 9: Plots of the absolute values of the two components of a solution from the class in Example 6.9, showing a bright, respectively dark soliton, a breather and a periodic rogue wave. We chose $\alpha=1$, $A_{1}=1/2$, $A_{2}=\sqrt{3}/2$, $\gamma=1$ (so that $r=-1/2+\mathrm{i}/2$), $c_{01}=c_{02}=c_{12}=c_{21}=c_{22}=\tilde{c}_{12}=1$, $c_{11}=\tilde{c}_{21}=-1/2+3\mathrm{i}/2$.
$\Box$
Example 6.10.
Let $\Gamma$ be the $2\times 2$ lower triangular Jordan block with eigenvalue $\gamma$. Then
$R=\pm\left(\begin{array}[]{cc}r&0\\ -(1-\mathrm{i}\,\alpha^{2}\beta\gamma)(4\gamma^{3}\,r)^{-1}&r\end{array}\right)\,,\qquad r=\frac{1}{2\gamma}\sqrt{1-\alpha^{2}\gamma\,(\gamma+2\mathrm{i}\,\beta)}\,.$ (6.42)
Let $r\neq 0$. We obtain
$\chi_{1}$ $=$ $e^{\frac{1}{2}\mathrm{i}\,\varphi}\left(\begin{array}[]{c}c_{11}\,e^{\psi}+c_{21}\,e^{-\psi}\\ (4\gamma^{3}r)^{-1}\Big(\big(-c_{11}\,\rho+4c_{12}\gamma^{3}r\big)\,e^{\psi}+\big(c_{21}\,\rho+4c_{22}\gamma^{3}r\big)\,e^{-\psi}\Big)\end{array}\right)\,$,
$\chi_{2}$ $=$ $e^{\mathrm{i}\,\varphi}\,e^{\frac{1}{2}(\gamma^{-1}x+\gamma\,t)}\left(\begin{array}[]{c}c_{01}\\ c_{02}-\frac{1}{2}c_{01}\gamma^{-1}(\gamma^{-1}x-\gamma\,t)\end{array}\right)\,$,
$\Xi$ $=$ $\frac{\|A\|}{8\gamma^{4}r}\,e^{\frac{\mathrm{i}}{2}\varphi}\left(\begin{array}[]{c}-4\gamma^{3}r\,\big(c_{11}\,b_{-}\,e^{\psi}+c_{21}\,b_{+}\,e^{-\psi}\big)\\ \big(c_{11}(b_{-}\,\rho-2\gamma\,d_{-})-4c_{12}b_{-}\gamma^{3}r\big)\,e^{\psi}-\big(c_{21}(b_{+}\,\rho-2\gamma\,d_{+})+4c_{22}b_{+}\gamma^{3}r\big)\,e^{-\psi}\end{array}\right)\,$,
where
$\rho=(1-\mathrm{i}\,\alpha^{2}\beta\gamma)\,x+\alpha\gamma^{2}(\beta-\mathrm{i}\,\gamma)\,t\,,\qquad\psi=r\,\big(x-\mathrm{i}\,\frac{\gamma}{\alpha}t\big)\,$,
$b_{\pm}=1\pm 2\gamma\,r+\mathrm{i}\,\alpha\,\gamma\,,\qquad d_{\pm}=1\pm 2\gamma\,r-\mathrm{i}\,\alpha^{2}\beta\,\gamma\,.$
It is now straightforward to elaborate the equations determining $\Omega$ and then the generated solution (6.21) of the two-component vector FL equation. If $\mathrm{Re}(\gamma)\neq 0$, then $\Omega$ is completely determined by the Lyapunov equation and given by Example A.2. If $\mathrm{Re}(\gamma)=0$, we have to employ the differential equations for $\Omega$ in addition, a more tedious task. We skip the details. $\Box$
Example 6.11.
Again, let $\Gamma$ be the $2\times 2$ lower triangular Jordan block with eigenvalue $\gamma$, so that $R$ and $r$ are given by (6.42). But now we consider the double root case $r=0$. Then the expression on the right hand side of (6.13) is nilpotent and does not possess a square root, and $\gamma$ is fixed by (6.36). The general solution of the linear system is now given by $\tilde{\chi}_{i}=(\tilde{\chi}_{i1},\tilde{\chi}_{i2})^{T}$, $i=1,2$, with
$\tilde{\chi}_{11}$ $=$ $a_{1}\,\rho+a_{0}\qquad\mbox{where}\quad\rho=x-\frac{\mathrm{i}}{2\gamma}\,t\,$,
$\tilde{\chi}_{12}$ $=$ $-\frac{(1-\mathrm{i}\,\alpha^{2}\beta\gamma)}{12\gamma^{3}}(a_{1}\,\rho^{3}+3a_{0}\,\rho^{2})+c_{2}\rho+c_{1}+\frac{6\,\mathrm{i}\,a_{1}}{12\gamma^{2}}\,t\,$,
$\tilde{\chi}_{21}$ $=$ $b_{1}\,e^{\frac{1}{2}(\gamma^{-1}x+\gamma\,t)}\,,\qquad\tilde{\chi}_{22}=\big(b_{2}-\frac{1}{2}b_{1}(\gamma^{-2}x-t)\big)\,e^{\frac{1}{2}(\gamma^{-1}x+\gamma\,t)}\,$,
with constants $a_{0},a_{1},b_{1},b_{2},c_{1},c_{2}$. Then we obtain the following components of $\Xi$,
$\Xi_{1}$ $=$ $-\frac{\|A\|}{2\gamma}\,e^{\mathrm{i}\,\varphi}\,\Big(a_{1}\,(\rho-2\gamma)+a_{0}\Big)\,$,
$\Xi_{2}$ $=$ $\|A\|\,e^{\mathrm{i}\,\varphi}\,\Big(\frac{a_{1}}{24\,\gamma^{4}}(1-\mathrm{i}\,\alpha^{2}\beta\gamma)\,\rho^{3}+\frac{a_{0}-2a_{1}\gamma}{8\gamma^{4}}(1-\mathrm{i}\,\alpha^{2}\beta\gamma)\,\rho^{2}$
$-\frac{1}{2\gamma^{3}}\big(a_{0}(1-\mathrm{i}\,\alpha^{2}\beta\gamma)-a_{1}\gamma+c_{2}\gamma^{2}\big)\,\rho+\frac{1}{4\gamma^{3}}\big(-\mathrm{i}\,a_{1}\,t+2a_{0}\gamma-2c_{1}\gamma^{2}+4c_{2}\gamma^{3}\big)\,.$
If $\alpha^{-1}>\|A\|^{2}$, then $\mathrm{Re}(\gamma)\neq 0$ (cf. Example 6.8), and $\Omega$ is completely determined by the Lyapunov equation alone,
$\Omega_{11}$ $=$ $\frac{1}{2\,\mathrm{Re}(\gamma)}\Big(\frac{\gamma^{\ast}}{\alpha^{2}\|A\|^{4}}|\Xi_{1}|^{2}+|\tilde{\chi}_{11}|^{2}+|\tilde{\chi}_{21}|^{2}\Big)\,$,
$\Omega_{12}$ $=$ $\frac{1}{4\,\mathrm{Re}(\gamma)^{2}\alpha^{2}\|A\|^{4}}\Big(\gamma\,|\Xi_{1}|^{2}+2\,\mathrm{Re}(\gamma)\,\gamma^{\ast}\,\Xi_{1}\,\Xi_{2}^{\ast}$
$-\alpha^{2}\|A\|^{4}\big(|\tilde{\chi}_{11}|^{2}+|\tilde{\chi}_{21}|^{2}-2\,\mathrm{Re}(\gamma)\,(\tilde{\chi}_{11}\tilde{\chi}_{12}^{\ast}+\tilde{\chi}_{21}\tilde{\chi}_{22}^{\ast})\big)\Big)\,$,
$\Omega_{21}$ $=$ $\frac{1}{4\,\mathrm{Re}(\gamma)^{2}\alpha^{2}\|A\|^{4}}\Big(-\gamma^{\ast}\,|\Xi_{1}|^{2}+2\,\mathrm{Re}(\gamma)\,\gamma^{\ast}\,\Xi_{1}^{\ast}\,\Xi_{2}$
$-\alpha^{2}\|A\|^{4}\big(|\tilde{\chi}_{11}|^{2}+|\tilde{\chi}_{21}|^{2}-2\,\mathrm{Re}(\gamma)\,(\tilde{\chi}_{11}^{\ast}\tilde{\chi}_{12}+\tilde{\chi}_{21}^{\ast}\tilde{\chi}_{22})\big)\Big)\,$,
$\Omega_{22}$ $=$ $\frac{1}{8\,\mathrm{Re}(\gamma)^{3}\alpha^{2}\|A\|^{4}}\Big(-2\mathrm{i}\,\mathrm{Im}(\gamma)\,|\Xi_{1}|^{2}+4\mathrm{Re}(\gamma)^{2}\,\gamma^{\ast}|\Xi_{2}|^{2}+4\mathrm{i}\,\mathrm{Re}(\gamma)\,\mathrm{Im}(\gamma\,\Xi_{1}^{\ast}\,\Xi_{2})$
$+2\alpha^{2}\|A\|^{4}\big(|\tilde{\chi}_{11}|^{2}+|\tilde{\chi}_{21}|^{2}+2\mathrm{Re}(\gamma)^{2}\,(|\tilde{\chi}_{12}|^{2}+|\tilde{\chi}_{22}|^{2})-2\,\mathrm{Re}(\gamma)\,\mathrm{Re}(\tilde{\chi}_{11}\tilde{\chi}_{12}^{\ast}+\tilde{\chi}_{21}\tilde{\chi}_{22}^{\ast})\big)\Big)\,.$
Now (6.21) determines a solution of the two-component vector FL equation, which is a second order version of the solution in Example 6.8. It has already been derived in a different way in [12]. This solution is quasi-rational if and only if $\tilde{\chi}_{2}=0$, in which case it is equivalent to a solution of the scalar FL equation. $\Box$
Remark 6.12.
The solution set of a polynomial matrix equation, like (6.13), is in general more intricate than for polynomial equations over a field. Here we take a look at an example in our context. Let $\Gamma=\mathrm{diag}(\gamma_{1},\gamma_{2})$, where $\gamma_{1},\gamma_{2}$ have the form (6.36). In this case,
@@
$\chi_{1}=e^{\frac{\mathrm{i}}{2}\,\varphi}\left(\begin{array}[]{c}c_{11}+c_{12}\,(x-\frac{\mathrm{i}}{\alpha}\gamma_{1}t)\\ c_{21}+c_{22}\,(x-\frac{\mathrm{i}}{\alpha}\gamma_{2}t)\end{array}\right)\,.$
@@
6.2 The generic plane wave case $\alpha_{1}\neq\alpha_{2}$